# Stochastic heat flow by moments

### v1

@@ 2.1. Delta-Bose semigroup
To reiterate, the constant actually depends on $\theta$, but we omit the dependence. It is proven in [GQT21] that on $L^{2}(\mathbb{R}^{2[\![n]\!]})$, $\mathcal{Q}^{[\![n]\!]}(t)$ is the semigroup of a self-adjoint operator whose spectrum is bounded from above. Hence, $\mathcal{Q}^{[\![n]\!]}(t)$ is strongly continuous on $L^{2}(\mathbb{R}^{2[\![n]\!]})$, meaning $\mathcal{Q}^{[\![n]\!]}(t)f\to f$ in $L^{2}(\mathbb{R}^{2[\![n]\!]})$ as $t\to 0$, for every $f\in L^{2}(\mathbb{R}^{2[\![n]\!]})$.
Next we consider the moments of $W$. Recall that, in the proof of Proposition 1.9, we decompose a GBF $Z$ into $Z=1+W$. We would like to perform a similar decomposition for an SHF. To this end, note that $\mathrm{Dgm}[\![1]\!]=\emptyset$; see (2.11). Therefore, $\mathcal{Q}^{[\![1]\!]}(t-s)=\mathcal{P}^{[\![1]\!]}(t-s)=\mathsf{p}(t-s)$, and any SHF $Z$ has the expectation $\mathbf{E}[Z_{s,t}\,g\otimes g^{\prime}]=\langle g,\mathsf{p}(t-s)g^{\prime}\rangle$ given by the heat kernel. Accordingly, we decompose any SHF $Z$ into $Z=\mathsf{p}^{\scriptscriptstyle\star}+W$, where
$\mathsf{p}^{\scriptscriptstyle\star}\in C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},M_{+}(\mathbb{R}^{4})),\qquad\mathsf{p}^{\scriptscriptstyle\star}_{s,t}\,f:=\int_{\mathbb{R}^{4}}\mathrm{d}x\mathrm{d}x^{\prime}\,\mathsf{p}(t-s,x-x^{\prime})f(x,x^{\prime})\ .$ (2.17)
By Definition 1.2(d), for $n=1,\ldots,4$,
@@ 2.2. Chapman–Kolmogorov equation and polymer measure
Our main goal here is to upgrade the Chapman–Kolmogorov equation in Definition 1.2(b) to $L^{\overline{n}}$ convergence and to allow any $L^{2}$-applicable mollifier. The result is stated in Corollary 1.7(b). In the process of doing so, we also obtained the finite-dimensional distributions of the polymer measure associated with $Z$. As mentioned previously, the polymer measure has been constructed in [CM24] as a limit of the noise-mollified SHE. Here we work with $Z$.
@@
for $\text{for every }r<\infty,\text{ there exists an }R<\infty\text{ such that }D\cap C_{R}(\mathbb{R}^{2})\text{ is dense in }C_{r}(\mathbb{R}^{2})\ ,$ (2.21a)
for $\text{for every }R<\infty,\text{ there exists a nonnegative }g\in D\text{ such that }\inf\{g(x)|\,|x|<R\}>0\ .$ (2.21b)
@@
$d(\mu,\nu):=\sum_{i_{1}=1}^{\infty}\cdots\sum_{i_{K}=1}^{\infty}2^{-i_{1}-\cdots-i_{K}}\wedge\big{|}(\mu-\nu)g_{D,i_{1}}\otimes\cdots\otimes g_{D,i_{K}}\big{|}\ ,\quad\mu,\nu\in M_{+}(\mathbb{R}^{2K})\ .$ (2.22)
Given (2.21), using the Banach–Alogulu Theorem and [Rud91, 3.8(a)], one can check that $d$ metrizes the vague topology on $M_{+}(\mathbb{R}^{2K})$, and using the Riesz–Markov–Kakutani representation theorem, see [Rud86, Theorems 2.14 and 2.17(b)], one can check that $d$ is complete.
To study the Chapman–Kolmogorov equation, similarly to [CM24, Sections 2.2–2.3], we start by defining certain products on $M_{+}(\mathbb{R}^{4})$.
@@
Let $\mu_{1},\ldots,\mu_{K-1}\in M_{+}(\mathbb{R}^{4})$ and let $u^{2},\ldots,u^{K-1}\in L^{2}(\mathbb{R}^{2})$ be nonnegative. We call $(\mu_{1}\circ_{u^{2}}\mu_{2}\cdots\circ_{u^{K-1}}\mu_{K-1})$ well-defined, if, for every $f\in C_{\mathrm{c}}(\mathbb{R}^{2[\![K]\!]})$, the integral
$\begin{split}\big{(}&\mu_{1}\circ_{u^{2}}\mu_{2}\cdots\circ_{u^{K-1}}\mu_{K-1}\big{)}f\\ &:=\int_{\mathbb{R}^{2K}}\mu_{1}(\mathrm{d}x_{1}\otimes\mathrm{d}x_{2})\prod_{k=2}^{K-1}u^{k}(x_{k+1}-x_{k})\mu_{k}(\mathrm{d}x_{k}\otimes\mathrm{d}x_{k+1})\cdot f(x_{1},\ldots,x_{K})\end{split}$ (2.23)
converges absolutely. In this case, by the Riesz–Markov–Kakutani theorem, the functional (2.23) defines an element of $M_{+}(\mathbb{R}^{2[\![K]\!]})$.
Next, take any $I:=\{k_{1}<\ldots<k_{|I|}\}\subset\{2,\ldots,K-1\}$. Call $(\mu_{1}\operatorname*{\circ}_{u^{2}}\cdots\operatorname*{\bullet}_{u^{k_{1}}}\mu_{k_{1}}\cdots\operatorname*{\bullet}_{u^{k_{2}}}\mu_{k_{2}}\cdots\mu_{K-1})$ well-defined, if, for every $f\in C_{\mathrm{c}}(\mathbb{R}^{2([\![K]\!]\setminus I)})$, the integral
$\begin{split}\big{(}&\mu_{1}\circ_{u^{2}}\cdots\bullet_{u^{k_{1}}}\mu_{k_{1}}\cdots\bullet_{u^{k_{2}}}\mu_{k_{2}}\cdots\mu_{K-1}\big{)}f\\ &:=\int_{\mathbb{R}^{2K}}\mu_{1}(\mathrm{d}x_{1}\otimes\mathrm{d}x_{2})\prod_{k=2}^{K-1}u^{k}(x_{k+1}-x_{k})\mu_{k}(\mathrm{d}x_{k}\otimes\mathrm{d}x_{k+1})\cdot f(\{x_{k}\}_{k\in[\![K]\!]\setminus I})\end{split}$ (2.24)
converges absolutely. In this case, by the Riesz–Markov–Kakutani theorem, the functional (2.24) defines an element of $M_{+}(\mathbb{R}^{2([\![K]\!]\setminus I)})$, which is the marginal of $\mu_{1}\circ_{u^{2}}\cdots\circ_{u^{K-1}}\mu_{K-1}$.
We now construct these products for $Z$. Since the results in this section will be used in the proof of Theorem 1.3, we make minimal assumptions on $Z$. The assumptions in the next lemma are satisfied by any SHF($\theta$) for $Z_{k}:=Z_{t_{k},t_{k+1}}$.
@@
Let $\overline{n}\in\mathbb{Z}_{>0}$, $T:=\{t_{1}<\ldots<t_{|T|}\}$, and $Z_{1},\ldots,Z_{|T|-1}$ be independent $M_{+}(\mathbb{R}^{4})$-valued random variables such that, for $n=1,\ldots,\overline{n}$, $k=1,\ldots,|T|-1$, and $g_{i},g^{\prime}_{i}\in L^{2}(\mathbb{R}^{2})$,
@@
Take nonnegative $u^{2},\ldots,u^{|T|-1}\in L^{2}(\mathbb{R}^{2})$ and $S=\{t_{i_{1}}<\ldots<t_{i_{|S|}}\}\subset T\setminus\{t_{1},t_{|T|-1}\}$ and let
$PM_{T,\vec{u}}$ $:=Z_{1}\circ_{u^{2}}Z_{2}\circ_{u^{3}}\cdots\circ_{u^{|T|-1}}Z_{|T|-1}\ ,$ (2.26)
$PM^{S}_{T,\vec{u}}$ $:=Z_{1}\circ_{u^{2}}\cdots Z_{i_{1}}\bullet_{u^{i_{1}}}\cdots\cdots Z_{i_{2}}\bullet_{u^{i_{2}}}\cdots\circ_{u^{|T|-1}}Z_{|T|-1}\ .$ (2.27)
Almost surely, $PM_{T,\vec{u}}$ and $PM^{S}_{T,\vec{u}}$ are well-defined.
@@
It suffices to show that, for every $g_{1},g_{|T|}\in C_{\mathrm{c}}(\mathbb{R}^{2})$ and $g_{2},\ldots,g_{|T|-1}\in C_{\mathrm{b}}(\mathbb{R}^{2})$,
$\mathbf{E}\big{[}(PM_{T,\vec{u}}|g_{1}|\otimes\cdots\otimes|g_{|T|}|)^{\overline{n}}\,\big{]}<\infty\ .$ (2.28)
Once this is done, since $D$ is countable, almost surely the condition holds: $PM_{T,\vec{u}}|g_{1}|\otimes\cdots\otimes|g_{|T|}|<\infty$ for all $g_{1},\ldots,g_{|T|}\in D$. This condition together with (2.21b) implies that $PM_{T,\vec{u}}|f|<\infty$ for all $f\in C_{\mathrm{c}}(\mathbb{R}^{2T})$. This shows that $PM_{T,\vec{u}}$ is almost surely well-defined. A similarly argument applies to $PM^{S}_{T,\vec{u}}$ by taking $g_{k}\in D$ for $t_{k}\in T\setminus S$ and $g_{k}=1$ for $t_{k}\in S$. To prove (2.28), use the independence of the $Z_{k}$s and (2.25) to write the expectation as
@@
Under the same setup as Lemma 2.3, let $\{u^{k}_{\ell}\}_{\ell}$, $k=2,\ldots,|T|-1$, be $L^{2}$-applicable mollifiers (Definition 1.1), let $\vec{u}_{\vec{\ell}}:=(u^{k}_{\ell_{k}})_{k=2,\ldots,|T|-1}$, and assume further $\overline{n}\in 2\mathbb{Z}_{>0}$. Then, there exist $(Z_{1},\ldots,Z_{|T|-1})$-measurable random variables $PM_{T}\in M_{+}(\mathbb{R}^{2T})$ and $PM^{S}_{T}\in M_{+}(\mathbb{R}^{2T\setminus S})$ such that, for every $f\in C_{\mathrm{c}}(\mathbb{R}^{2T})$ and $h\in C_{\mathrm{c}}(\mathbb{R}^{2T\setminus S})$,
$\mathbf{E}\big{[}\big{|}\big{(}PM_{T,\vec{u}_{\vec{\ell}}}-PM_{T}\big{)}f\big{|}^{\overline{n}}\big{]}\ ,\ \mathbf{E}\big{[}\big{|}\big{(}PM^{S}_{T,\vec{u}_{\vec{\ell}}}-PM^{S}_{T}\big{)}h\big{|}^{\overline{n}}\big{]}\longrightarrow 0\ \text{ as }\vec{\ell}\to\vec{\infty}\ ,$ (2.30)
and the limits $PM_{T}$ and $PM^{S}_{T}$ do not depend on the mollifiers.
@@
Let $\{v^{1}_{\ell}\},\ldots,\{v^{k}_{\ell}\}$ also be $L^{2}$-applicable mollifiers. Given (2.21) and that $d$ is complete, it suffices to prove that, for every $g_{1},g_{|T|}\in C_{\mathrm{c}}(\mathbb{R}^{2})$ and every $g_{2},\ldots,g_{|T|-1}\in C_{\mathrm{b}}(\mathbb{R}^{2})$,
$\mathbf{E}\big{[}\big{|}(PM_{T,\vec{u}_{\vec{\ell}}}-PM_{T,\vec{v}_{\vec{\ell}^{\prime}}})\,g_{1}\otimes\cdots\otimes g_{|T|}\big{|}^{\overline{n}}\big{]}\longrightarrow 0\quad\text{ as }(\vec{\ell},\vec{\ell}^{\prime})\to(\vec{\infty},\vec{\infty})\ .$ (2.31)
@@
$\mathbf{E}\Big{[}\big{(}(PM_{T,\vec{u}_{\vec{\ell}}}-PM_{T,\vec{v}_{\vec{\ell}^{\prime}}})\,g_{1}\otimes\cdots\otimes g_{|T|}\big{)}^{\overline{n}}\Big{]}=\sum_{j=0}^{\overline{n}}\binom{\overline{n}}{j}(-1)^{j}A_{j,\vec{\ell},\vec{\ell}^{\prime}}\ ,$ (2.32)
where $A_{j,\vec{\ell},\vec{\ell}^{\prime}}:=\mathbf{E}[(PM_{T,\vec{u}_{\ell}}\,g_{1}\otimes\cdots\otimes g_{|T|})^{\overline{n}-j}(PM_{T,\vec{v}_{\ell^{\prime}}}\,g_{1}\otimes\cdots\otimes g_{|T|})^{j}]$. Use the independence of the $Z_{k}$s and (2.25) to evaluate the last expectation. Doing so gives
$A_{j,\vec{\ell},\vec{\ell}^{\prime}}=\Big{\langle}g_{1}^{\otimes\overline{n}},\mathcal{Q}^{[\![\overline{n}]\!]}(t_{2}-t_{1})\prod_{k=2}^{|T|-1}g_{k}^{\otimes\overline{n}}\,\big{(}u^{k,\star}_{\ell_{k}}{}^{\otimes(\overline{n}-j)}\otimes v^{k,\star}_{\ell_{k}^{\prime}}{}^{\otimes j}\big{)}\,\mathcal{Q}^{[\![\overline{n}]\!]}(t_{k+1}-t_{k})\cdot g_{|T|}^{\otimes\overline{n}}\Big{\rangle}\ .$ (2.33)
Because $\{u^{i}_{\ell}\}$ and $\{v^{i}_{\ell}\}$ are $L^{2}$-applicable mollifiers, as $(\vec{\ell},\vec{\ell}^{\prime})\to(\vec{\infty},\vec{\infty})$, the term $A_{j,\vec{\ell},\vec{\ell}^{\prime}}$ converges to $\langle g_{1}^{\otimes\overline{n}},\mathcal{Q}^{[\![\overline{n}]\!]}(t_{2}-t_{1})\prod_{k=2}^{|T|-1}g_{k}^{\otimes\overline{n}}\mathcal{Q}^{[\![\overline{n}]\!]}(t_{k+1}-t_{k})\cdot g_{|T|}^{\otimes\overline{n}}\rangle$. Since this limit is independent of $j$, the limit of (2.32) is zero, which verifies (2.31). ∎
Now take any $Z$ that satisfies Definition 1.2(b), (c), and (d) for an $\overline{n}\in 2\mathbb{Z}_{>0}$ and apply Lemma 2.3 and Proposition 2.4 with $Z_{k}:=Z_{t_{k},t_{k+1}}$. The resulting limit, still written as $PM_{T}\in M_{+}(\mathbb{R}^{2T})$, is the finite dimensional distribution of the polymer measure associated with an SHF($\theta$). Once Theorem 1.3 and Proposition 1.4 are proven, the polymer measure exists and is unique in law. Similarly, $PM^{S}_{T}$ is the marginal of $PM_{T}$ on $\mathbb{R}^{2(T\setminus S)}$. Combining Proposition 2.4 and Definition 1.2(b) gives the consistent properties.
@@
    $PM^{S}_{T}=PM_{T\setminus S}$ almost surely.
@@
    For any $L^{2}$-applicable mollifiers $\{u^{k}_{\ell}\}_{\ell}$, $k=2,\ldots,|T|-1$, and any $f\in C_{\mathrm{c}}(\mathbb{R}^{4})$,
    $\mathbf{E}\big{[}\big{|}\big{(}Z_{t_{1},t_{2}}\bullet_{u^{2}_{\ell_{2}}}\cdots\bullet_{u^{|T|-1}{}_{\ell_{|T|-1}}}Z_{t_{|T|-1},t_{|T|}}-Z_{t_{1},t_{|T|}}\big{)}f\big{|}^{\overline{n}}\big{]}\longrightarrow 0\ ,\qquad\text{as }\vec{\ell}\to\vec{\infty}\ .$ (2.34)
3. Uniqueness, proof of Theorem 1.3
We start the proof with a few reductions. Let $Z$ and $\widetilde{Z}$ be SHF($\theta$). By Definition 1.2(a) and (c) and Corollary 2.5(b), it suffices to prove that, for every $s<t$, $Z_{s,t}$ and $\widetilde{Z}_{s,t}$ have the same law. Take $(s,t)=(0,1)$ to simplify notation. By a standard criterion, see [Kal17, Theorem 2.2] for example, it suffices to prove that, for every $h\in C_{\mathrm{b}}(\mathbb{R})$ and $f\in C_{\mathrm{c}}(\mathbb{R}^{4})$, $\mathbf{E}[h(Z_{0,1}f)]-\mathbf{E}[h(\widetilde{Z}_{0,1}f)]=0$. Via an approximation argument, we can assume that $h$ is $C^{4}$ with bounded derivatives and that $f\in C_{\mathrm{c}}(\mathbb{R}^{2})^{\otimes 2}$, more explicitly

### v3

@@ 2.1. Delta-Bose semigroup
To reiterate, the constant actually depends on $\theta$, but we omit the dependence. It is proven in [GQT21] that on $L^{2}(\mathbb{R}^{2[\![n]\!]})$, $\mathcal{Q}^{[\![n]\!]}(t)$ is the semigroup generated by a self-adjoint operator whose spectrum is bounded from above. Hence, $\mathcal{Q}^{[\![n]\!]}(t)$ is strongly continuous on $L^{2}(\mathbb{R}^{2[\![n]\!]})$, meaning $\mathcal{Q}^{[\![n]\!]}(t)f\to f$ in $L^{2}(\mathbb{R}^{2[\![n]\!]})$ as $t\to 0$, for every $f\in L^{2}(\mathbb{R}^{2[\![n]\!]})$.
Next, we consider the moments of the increment $W:=Z-\mathbf{E}Z$. First, since $\mathrm{Dgm}[\![1]\!]=\emptyset$ (see (2.11)), $\mathcal{Q}^{[\![1]\!]}(t-s)=\mathcal{P}^{[\![1]\!]}(t-s)=\mathsf{p}(t-s)$, and any SHF $Z$ has its expectation $\mathbf{E}[Z_{s,t}\,g\otimes g^{\prime}]=\langle g,\mathsf{p}(t-s)g^{\prime}\rangle$ given by the heat kernel. Hence, letting
$\mathsf{p}^{\scriptscriptstyle\star}\in C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathcal{M}_{+}(\mathbb{R}^{4})),\qquad\mathsf{p}^{\scriptscriptstyle\star}_{s,t}\,f:=\int_{\mathbb{R}^{4}}\mathrm{d}x\mathrm{d}x^{\prime}\,\mathsf{p}(t-s,x-x^{\prime})f(x,x^{\prime})\ ,$ (2.17)
we have $W=Z-\mathsf{p}^{\scriptscriptstyle\star}$. By Definition 1.2(d), for $n=1,\ldots,4$,
@@ 2.2. Chapman–Kolmogorov equation and polymer measure
Our main goal here is to upgrade the Chapman–Kolmogorov equation in Definition 1.2(b) to $L^{\overline{n}}$ convergence and to allow any $L^{2}$-applicable mollifier. The result is stated in Corollary 1.8(b). In the process of doing so, we also obtained the finite-dimensional distributions of the polymer measure associated with $Z$. As mentioned previously, the polymer measure has been constructed in [CM24] as a limit of the noise-mollified SHE.
@@
[…]
@@
$d(\mu,\nu):=\sum_{i_{1}=1}^{\infty}\cdots\sum_{i_{K}=1}^{\infty}2^{-i_{1}-\cdots-i_{K}}\wedge\big|(\mu-\nu)g_{D,i_{1}}\otimes\cdots\otimes g_{D,i_{K}}\big|\ ,\quad\mu,\nu\in\mathcal{M}_{+}(\mathbb{R}^{2K})\ .$ (2.22)
Given (2.21), using the Banach–Alogulu Theorem and [Rud91, 3.8(a)], one can check that $d$ metrizes the vague topology on $\mathcal{M}_{+}(\mathbb{R}^{2K})$, and using the Riesz–Markov–Kakutani representation theorem, see [Rud86, Theorems 2.14], one can check that $d$ is complete.
To study the Chapman–Kolmogorov equation, similarly to [CM24, Sections 2.2–2.3], we start by defining certain products in $\mathcal{M}_{+}(\mathbb{R}^{4})$.
@@
Let $\mu_{1},\ldots,\mu_{K}\in\mathcal{M}_{+}(\mathbb{R}^{4})$ and let $u^{1},\ldots,u^{K-1}\in L^{2}(\mathbb{R}^{2})\cap C(\mathbb{R}^{2})$ be nonnegative. We call $(\mu_{1}\circ_{u^{1}}\cdots\circ_{u^{K-1}}\mu_{K})$ well-defined, if, for every $f\in C_{\mathrm{c}}(\mathbb{R}^{2\{0,\ldots,K\}})$, the integral
$\begin{split}\big(&\mu_{1}\circ_{u^{1}}\cdots\circ_{u^{K-1}}\mu_{K}\big)f\\ &:=\int_{\mathbb{R}^{4K}}\mu_{1}(\mathrm{d}x_{0}\otimes\mathrm{d}x_{1})\prod_{k=2}^{K}u^{k-1}(x_{k-\frac{1}{2}}-x_{k-1})\mu_{k}(\mathrm{d}x_{k-\frac{1}{2}}\otimes\mathrm{d}x_{k})\cdot f(x_{0},\ldots,x_{K})\end{split}$ (2.23)
converges absolutely. In this case, by the Riesz–Markov–Kakutani theorem, the functional (2.23) defines an element of $\mathcal{M}_{+}(\mathbb{R}^{2K+2})$.
Next, take any $I:=\{k_{1}<\ldots<k_{|I|}\}\subset\{1,\ldots,K-1\}$. Call $(\mu_{1}\circ_{u^{1}}\cdots\bullet_{u^{k_{1}}}\mu_{k_{1}}\cdots\bullet_{u^{k_{2}}}\mu_{k_{2}}\cdots\mu_{K})$ well-defined, if, for every $f\in C_{\mathrm{c}}(\mathbb{R}^{2(\{0,\ldots,K\}\setminus I)})$, the integral
$\begin{split}\big(&\mu_{1}\circ_{u^{1}}\cdots\mu_{k_{1}}\bullet_{u^{k_{1}}}\cdots\mu_{k_{2}}\bullet_{u^{k_{2}}}\cdots\mu_{K}\big)f\\ &:=\int_{\mathbb{R}^{4K}}\mu_{1}(\mathrm{d}x_{0}\otimes\mathrm{d}x_{1})\prod_{k=2}^{K}u^{k-1}(x_{k-\frac{1}{2}}-x_{k-1})\mu_{k}(\mathrm{d}x_{k-\frac{1}{2}}\otimes\mathrm{d}x_{k})\cdot f(\{x_{k}\}_{k\in\{0,\ldots,K\}\setminus I})\end{split}$ (2.24)
converges absolutely. In this case, by the Riesz–Markov–Kakutani theorem, the functional (2.24) defines an element of $\mathcal{M}_{+}(\mathbb{R}^{2(\{0,\ldots,K\}\setminus I)})$, which is the marginal of $\mu_{1}\circ_{u^{1}}\cdots\circ_{u^{K-1}}\mu_{K}$.
We now construct these products for $Z$. Since the results in this section will be used in the proof of Theorem 1.3, we make minimal assumptions on $Z$. The assumptions in the next lemma are satisfied by any SHF($\theta$) for $Z_{k}:=Z_{t_{k-1},t_{k}}$.
@@
Take any $\overline{n}\in\mathbb{Z}_{>0}$, $U=\{t_{1}<\ldots<t_{|U|}\}$, and $t_{0}\in(-\infty,t_{1})$, let $U^{\prime}:=\{t_{0}<t_{1}<\ldots<t_{|U|}\}$, and let $Z_{1},\ldots,Z_{|U|}$ be independent $\mathcal{M}_{+}(\mathbb{R}^{4})$-valued random variables such that, for $n=1,\ldots,\overline{n}$, $k=1,\ldots,|U|$, and $g_{i},g^{\prime}_{i}\in L^{2}(\mathbb{R}^{2})$,
@@
Take nonnegative $u^{1},\ldots,u^{|U|-1}\in L^{2}(\mathbb{R}^{2})\cap C(\mathbb{R}^{2})$ and $V=\{t_{i_{1}}<\ldots<t_{i_{|V|}}\}\subset\{t_{1},\ldots,t_{|U|-1}\}$, and let
$M_{U^{\prime},\vec{u}}$ $:=Z_{1}\circ_{u^{1}}Z_{2}\circ_{u^{2}}\cdots Z_{|U|}\ ,$ (2.26)
$M^{V}_{U^{\prime},\vec{u}}$ $:=Z_{1}\circ_{u^{1}}\cdots Z_{i_{1}}\bullet_{u^{i_{1}}}\cdots Z_{i_{2}}\bullet_{u^{i_{2}}}\cdots Z_{|U|}\ .$ (2.27)
Almost surely, $M_{U^{\prime},\vec{u}}$ and $M^{V}_{U^{\prime},\vec{u}}$ are well-defined.
@@
It suffices to show that, for every $g_{0},g_{|U|}\in C_{\mathrm{c}}(\mathbb{R}^{2})$ and $g_{1},\ldots,g_{|U|-1}\in C_{\mathrm{b}}(\mathbb{R}^{2})$,
$\mathbf{E}\big[(M_{U^{\prime},\vec{u}}\,|g_{0}|\otimes\cdots\otimes|g_{|U|}|)^{\overline{n}}\,\big]<\infty\ .$ (2.28)
Once this is done, since $D$ is countable, almost surely the condition holds: $M_{U,\vec{u}}|g_{0}|\otimes\cdots\otimes|g_{|U|}|<\infty$ for all $g_{0},\ldots,g_{|U|}\in D$. This condition together with (2.21b) implies that $M_{U^{\prime},\vec{u}}|f|<\infty$ for all $f\in C_{\mathrm{c}}(\mathbb{R}^{2U^{\prime}})$. This shows that $M_{U^{\prime},\vec{u}}$ is almost surely well-defined. A similarly argument applies to $M^{V}_{U^{\prime},\vec{u}}$ by taking $g_{k}\in D$ for $t_{k}\in U\setminus V$ and $g_{k}=1$ for $t_{k}\in V$. To prove (2.28), use the independence of the $Z_{k}$s and (2.25) to write the expectation as
@@
Under the same setup as Lemma 2.3, let $\{u^{k}_{\ell}\}_{\ell}$, $k=1,\ldots,|U|-1$, be $L^{2}$-applicable mollifiers (Definition 1.1), let $\vec{u}_{\vec{\ell}}:=(u^{k}_{\ell_{k}})_{k=1,\ldots,|U|-1}$, and assume further $\overline{n}\in 2\mathbb{Z}_{>0}$. Then, there exist $(Z_{1},\ldots,Z_{|U|})$-measurable random variables $M_{U^{\prime}}\in\mathcal{M}_{+}(\mathbb{R}^{2U^{\prime}})$ and $M^{V}_{U^{\prime}}\in\mathcal{M}_{+}(\mathbb{R}^{2U^{\prime}\setminus V})$ such that, for every $f\in C_{\mathrm{c}}(\mathbb{R}^{2U^{\prime}})$ and $h\in C_{\mathrm{c}}(\mathbb{R}^{2U^{\prime}\setminus V})$,
[…]
and the limits $M_{U^{\prime}}$ and $M^{V}_{U^{\prime}}$ do not depend on the mollifiers.
@@
Let $\{v^{1}_{\ell}\},\ldots,\{v^{k}_{\ell}\}$ also be $L^{2}$-applicable mollifiers. Given (2.21) and that $d$ is complete, it suffices to prove that, for every $g_{0},g_{|U|}\in C_{\mathrm{c}}(\mathbb{R}^{2})$ and every $g_{1},\ldots,g_{|U|-1}\in C_{\mathrm{b}}(\mathbb{R}^{2})$,
$\mathbf{E}\big[\big|(M_{U^{\prime},\vec{u}_{\vec{\ell}}}-M_{U^{\prime},\vec{v}_{\vec{\ell}^{\prime}}})\,g_{0}\otimes\cdots\otimes g_{|U|}\big|^{\overline{n}}\big]\longrightarrow 0\quad\text{ as }(\vec{\ell},\vec{\ell}^{\prime})\to(\vec{\infty},\vec{\infty})\ .$ (2.31)
@@
$\mathbf{E}\Big[\big((M_{U^{\prime},\vec{u}_{\vec{\ell}}}-M_{U^{\prime},\vec{v}_{\vec{\ell}^{\prime}}})\,g_{0}\otimes\cdots\otimes g_{|U|}\big)^{\overline{n}}\Big]=\sum_{j=0}^{\overline{n}}\binom{\overline{n}}{j}(-1)^{j}A_{j,\vec{\ell},\vec{\ell}^{\prime}}\ ,$ (2.32)
where $A_{j,\vec{\ell},\vec{\ell}^{\prime}}:=\mathbf{E}[(M_{U^{\prime},\vec{u}_{\ell}}\,g_{0}\otimes\cdots\otimes g_{|U|})^{\overline{n}-j}(M_{U^{\prime},\vec{v}_{\ell^{\prime}}}\,g_{0}\otimes\cdots\otimes g_{|U|})^{j}]$. Use the independence of the $Z_{k}$s and (2.25) to evaluate the last expectation. Doing so gives
$A_{j,\vec{\ell},\vec{\ell}^{\prime}}=\Big\langle g_{0}^{\otimes\overline{n}},\mathcal{Q}^{[\![\overline{n}]\!]}(t_{1}-t_{0})\prod_{k=1}^{|U|-1}g_{k}^{\otimes\overline{n}}\,\big(u^{k,\star}_{\ell_{k}}{}^{\otimes(\overline{n}-j)}\otimes v^{k,\star}_{\ell_{k}^{\prime}}{}^{\otimes j}\big)\,\mathcal{Q}^{[\![\overline{n}]\!]}(t_{k+1}-t_{k})\cdot g_{|U|}^{\otimes\overline{n}}\Big\rangle\ .$ (2.33)
[…]
Remark 2.5.
The martingale formulation of the products $\bullet$ and $\circ$
Now take any $Z$ that satisfies Definition 1.2(b), (c), and (d) for an $\overline{n}\in 2\mathbb{Z}_{>0}$ and apply Lemma 2.3 and Proposition 2.4 with $Z_{k}:=Z_{t_{k},t_{k+1}}$. The resulting limit, still written as $M_{U^{\prime}}\in\mathcal{M}_{+}(\mathbb{R}^{2U})$, is the finite dimensional distribution of the polymer measure associated with an SHF($\theta$). Once Theorem 1.3 and Proposition 1.5 are proven, the polymer measure exists and is unique in law. Similarly, $M^{V}_{U^{\prime}}$ is the marginal of $M_{U^{\prime}}$ on $\mathbb{R}^{2(U^{\prime}\setminus S)}$. Combining Proposition 2.4 and Definition 1.2(b) gives the consistent properties.
@@
    $M^{V}_{U^{\prime}}=M_{U^{\prime}\setminus V}$ almost surely.
@@
    For any $L^{2}$-applicable mollifiers $\{u^{k}_{\ell}\}_{\ell}$, $k=1,\ldots,|U|-1$, and any $f\in C_{\mathrm{c}}(\mathbb{R}^{4})$,
[…]
3. Axiomatic characterization, proof of Theorem 1.9
As mentioned after Theorem 1.9, it immediately implies Theorem 1.3.
We start the proof of Theorem 1.9 with a few reductions. Take any $G$, $Z$, and $\widetilde{Z}$ as in Theorem 1.9. By Definition 1.2(a) and (c) and Corollary 2.6(b), it suffices to prove that, for every $s<t\in G$, $Z_{s,t}$ and $\widetilde{Z}_{s,t}$ have the same law. To simplify notation, take $(s,t)=(0,1)$ with the assumption that $0,1\in G$. The proof works for any $s<t\in G$ without the assumption that $0,1\in G$. By a standard criterion, see [Kal17, Theorem 2.2] for example, it suffices to prove that, for every $h\in C_{\mathrm{b}}(\mathbb{R})$ and $f\in C_{\mathrm{c}}(\mathbb{R}^{4})$, $\mathbf{E}[h(Z_{0,1}f)]-\mathbf{E}[h(\widetilde{Z}_{0,1}f)]=0$. Via an approximation argument, we can assume that $h$ is $C^{4}$ with bounded derivatives and that $f\in C_{\mathrm{c}}(\mathbb{R}^{2})^{\otimes 2}$, more explicitly